### v1

# Samson Lasaulce, Giuseppe Valenzise, Zhuo He, and Merouane Debbah

Abstract
The development of wireless sensing technologies, using signals such as Wi-Fi, infrared, and RF to gather environmental data, has significantly advanced within Internet of Things (IoT) systems. Among these, Radio Frequency (RF) sensing stands out for its cost-effective and non-intrusive monitoring of human activities and environmental changes. However, traditional RF sensing methods face significant challenges, including noise, interference, incomplete data, and high deployment costs, which limit their effectiveness and scalability. This paper investigates the potential of Generative AI (GenAI) to overcome these limitations within the IoT ecosystem. We provide a comprehensive review of state-of-the-art GenAI techniques, focusing on their application to RF sensing problems. By generating high-quality synthetic data, enhancing signal quality, and integrating multi-modal data, GenAI offers robust solutions for RF environment reconstruction, localization, and imaging. Additionally, GenAI’s ability to generalize enables IoT devices to adapt to new environments and unseen tasks, improving their efficiency and performance. The main contributions of this article include a detailed analysis of the challenges in RF sensing, the presentation of innovative GenAI-based solutions, and the proposal of a unified framework for diverse RF sensing tasks. Through case studies, we demonstrate the effectiveness of integrating GenAI models, leading to advanced, scalable, and intelligent IoT systems.
Index Terms:
Generative AI, RF sensing, cross-modal estimation, multi-modal fusion, large language models.
I Introduction
With the development of the Internet of Things (IoT), many kinds of wireless sensing signals (e.g., Wi-Fi, Infrared images, visible images, Radio Frequency (RF) signal) are filling our living and working spaces nowadays. Recently, researchers have also utilized RF signals to capture events in the IoT environment (i.e., RF sensing). While RF signals are transmitted, reflected, blocked, and scattered by objects like walls, furniture, vehicles, and human bodies, it is possible to extract useful information, such as position, movement direction, speed, and vital signs of a human subject, from received RF signals. Unlike traditional hardware sensors, RF sensing provides users with low-cost and unobtrusive services. Furthermore, due to the broadcast nature of RF signals, RF sensing can be used not only to monitor multiple subjects, but also to capture changes in the environment over a large area [1].
Traditional RF sensing methods face several limitations in IoT systems, including noise and interference, which degrade signal quality and lead to inaccurate data interpretation. The presence of other electronic devices and RF sources causes further data loss or corruption. Incomplete data is common in scenarios with limited sensor deployment, and the high costs of deploying and maintaining extensive sensor networks make large-scale implementations expensive. Additionally, unstable environments cause signal weakening and multipath propagation, reducing reliability. These challenges necessitate advanced solutions like Generative AI (GenAI) to enhance the robustness, efficiency, and scalability of IoT systems.
GenAI refer to neural network models designed to generate new data similar to a given dataset, including Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), Autoregressive Models, flow-Based Models, Diffusion Models (DMs), and Transformer-based Large Language Models (LLMs). These techniques offer significant advantages in data-intensive applications by creating high-quality synthetic data, improving data quality through denoising, and filling in missing values. Generative AI is particularly effective in both cross-modal and multi-modal applications: integrating diverse data types into unified representations for better decision-making and translating information between modalities to enhance robustness. This capability supports innovative IoT applications, smart cities, healthcare, and autonomous systems, showcasing generative AI’s transformative potential.
Figure 1: This figure demonstrates an application scenario of RF sensing in IoT systems, highlighting the enhancements provided by generative AI techniques. These enhancements include RF data augmentation to complete missing observations and augment limited real-world data, cross-modality generation to infer missing data from available modalities, and multi-modal fusion to combine information from different data types with RF signals. The figure showcases the use of GANs, VAEs, DMs, and LLMs to address these challenges and improve the overall effectiveness of RF sensing in IoT applications.
GenAI’s ability to enhance data quality and integrate various data types makes it ideal for IoT applications, which require universality. With the advent of smarter devices, advanced sensors, and enhanced connectivity technologies like 5G and 6G, IoT systems can greatly benefit from GenAI. It extends conventional deep learning to manage diverse and unforeseen tasks with limited data and resources. GenAI’s generalization capability is crucial for IoT devices to adapt to new environments and tasks. Additionally, GenAI’s natural language processing enhances multi-modal sensing by integrating text, audio, and visual data, creating more comprehensive and intelligent IoT systems.
As shown in Fig. 1, we discuss the main challenges in RF sensing applications and explore how GenAI can address these issues using unimodal and multi-modal datasets, including reviewing the most relevant works and proposing feasible solutions for the potential use of GenAI. The main contributions of this article can be summarized as follows:
  • We provide a comprehensive review of state-of-the-art GenAI techniques in RF sensing, covering both uni-modal and multi-modal challenges. We highlight how GenAI models address key RF sensing issues, leveraging their statistical characteristics to model complex data distributions, handle missing data, and enhance signal processing.
  • We discuss the challenges in IoT systems posed by vast data volumes, numerous smart devices, task complexity, and the need for high generalization, particularly in RF sensing. We propose addressing these challenges using GenAI, including LLMs.
  • We propose a unified RF sensing framework for individual, multiple, and general tasks, and illustrate the effectiveness of integrating GenAI models with a case study.
II A short review of GenAI techniques
In this section, a short review of GenAI and their potential in RF sensing applications are provided (see Table I). GenAI is defined as AI models which are capable of generating new contents, data, or information from the learned distribution of the trained data. We provide a brief introduction to popular GenAI models and review their application to RF sensing, respectively.
GANs [2] generally consist of two neural networks interacting in the form of a zero-sum game, in which the generator $G$ aims to generate new data with the same statistics as the training dataset while the discriminator $D$ focuses on the identification of the true and fake data samples. In RF sensing applications, GANs offer the potential to generate realistic synthetic RF data, enhance signal quality, and enable robust anomaly detection. Their ability to model complex signal environments and facilitate data augmentation paves the way for more resilient and adaptive RF sensing systems.
VAEs learn an encoder and a decoder to embed data into a continuous latent space, where the decoder uses a parameterized likelihood function [3]. This approach reduces overfitting, making VAEs ideal for large datasets like radio map construction. In RF sensing, VAEs significantly enhance performance by denoising data and reconstructing incomplete information. By understanding the structure and distribution of RF signals, VAEs filter out noise and infer missing parts, creating comprehensive datasets even with partial sensor coverage or data loss. These capabilities make RF sensing systems more robust, adaptable, and effective, improving their performance and reliability in various IoT applications.
DMs use a Markov chain to add random noise to data (forward diffusion) and then denoise it to generate desired samples (reverse diffusion) [4]. The forward process transforms the data distribution into a simple prior (e.g., Gaussian noise), and new data is generated by denoising samples from this prior. Diffusion models enhance RF sensing in IoT by iteratively generating and denoising data, producing high-quality synthetic RF data and improving signal clarity. They effectively detect anomalies and capture fine variations in RF environments, and robustly reconstruct incomplete RF data, ensuring comprehensive datasets despite partial sensor coverage. These capabilities improve the precision, reliability, and robustness of RF sensing in IoT applications.
LLMs [5] are pre-trained on massive text datasets to learn contextual dependencies, and have achieved remarkable success in various applications, including question answering, language comprehension, reading comprehension, code creation, and both mathematical and common sense reasoning. In IoT systems, LLMs enhance RF sensing by incorporating natural language processing, enabling sensors to interpret and generate human language for smarter device communication. By involving the natural language modality, LLMs facilitate sophisticated multi-modal data analysis, combining RF data with textual and audio inputs for comprehensive insights. This capability improves anomaly detection, contextual understanding, and decision-making processes, making IoT systems more intelligent and adaptive.
TABLE I: The outline of issues encountered in RF sensing for IoT applications, the limitations of traditional AI techniques in addressing these challenges, and the potential of generative AI to enhance performance and efficiency in these contexts.
{tblr}
width = colspec = Q[200]Q[362]Q[378], cells = Thistle, valign = m, cells = BlueChalk, valign = m, column2 = Seashell, cell12 = fg=MineShaft, cell13 = fg=MineShaft, cell22 = fg=MineShaft, hlines, vline2-3 = -, Principal issues & Limitation of traditional deep learning techniques Potential of Generative AI
$\cdot$ Incomplete RF data;
$\cdot$ Limited training data in new environments. $\cdot$ Dependency on large and clean datasets;
$\cdot$ Lack of Built-in Mechanisms for Missing Data. $\cdot$ Data augmentation and synthesis;
$\cdot$ Learning data distributions;
$\cdot$ Generated plausible values from learned data distribution.
$\cdot$ Strong noise;
$\cdot$ Interference. $\cdot$ Trained on data with specific noise characteristics;
$\cdot$ Limited robustness to varying conditions;
$\cdot$ Tendency to overfit to noise. $\cdot$ Learn the underlying distribution of both signal and noise;
$\cdot$ Generalize to different noise conditions;
$\cdot$ Synthesizing training data with various noise and interference scenarios.
$\cdot$ Missing modalities;
$\cdot$ Unstable acquisition during test. $\cdot$ Require complete datasets for training and inference;
$\cdot$ Missing data result in degraded performance. $\cdot$ Integrate information from multiple modalities;
$\cdot$ Compensate for missing data in one modality using the available data from others.
$\cdot$ Fuse information in different modality;
$\cdot$ Low communication efficiency among Internet of Sensors $\cdot$ Static network structures;
$\cdot$ Difficult to exchange information among sensors;
$\cdot$ Difficult to adapt to different tasks and dynamic environment. $\cdot$ Provide potential for communication among sensors and intelligent edges;
$\cdot$ Provide convenience for dynamical environment adaptation.
III GenAI empowered uni-modal wireless sensing in IoT
The field of RF sensing involves utilizing radio frequency signals to detect and interpret physical phenomena, offering critical capabilities for applications such as environmental monitoring, health diagnostics, and security systems. Both traditional and modern RF sensing techniques face inherent challenges during the data acquisition and transmission stages. Among these challenges, high costs for data collection, significant interference during data acquisition, and bandwidth limitations for the data transmission are particularly prominent. Additionally, although RF signals are considered crucial information source in many IoT applications, they are extremely sparse due to the difficulty and high costs associated with setting up extensive observation points. To address these issues, we explore the emerging application of GenAI techniques to uni-modality RF sensing, building on their success in the computer vision domain, and provide visions for future directions.
Moreover, despite the advantages of using RF signals in IoT systems, challenges arise due to the rapid increase in IoT devices, which complicates efficient signal acquisition and data collection. Precise synchronization, robust interference handling, and managing large data volumes are particularly challenging. Consequently, RF signals often suffer from missed observations and exhibit a sparse structure, creating obstacles for effective downstream tasks. Generative AI addresses these challenges by using models like GANs, VAEs, and DMs to reconstruct missing data, infer information for localization, and generate representative samples for downstream tasks (see Fig. 2). The advantages of GenAI in the unimodal case can be discussed in three main aspects.
III-A Data imputation
A specific issue within RF sensing is the handling of missing Received Signal Strength (RSS) readings, which can lead to low accuracy and reliability in various applications, such as localization and environment monitoring. In these cases, the missing data needs to be filled in based on the observed data to minimize the impact. In the context of filling missing data, the correlation between the missing data and observed data is crucial. However, this correlation varies and traditional neural networks face challenges in handling this variability, as they often assume fixed spatial relationships or rely on predefined spatial structures. To address this issue, GenAI models, such as Transformers, have a great potential in learning complex dependencies and varying impacts of different observations, in which attention mechanisms can explicitly model the varying impact of each observation. Additionally, GenAI can create embeddings that capture broader context and relationships within the data, which can generate more accurate missing data. For instance, Wang et al.[6] leverages the Bidirectional Encoder Representations from Transformers (BERT) model, known for its strong representation learning ability and generalization capability in natural language processing, to input collected RSS data and predict missing signals in the radio map. In the case of RF sensing, BERT can be adapted to understand the contextual relationships within RSS data, much like it understands word relationships in a sentence. This capability allows BERT to model complex dependencies and patterns in the RSS data, making it highly effective at inferring missing values. By accurately predicting these missing signals, BERT helps to reduce the workload of collecting RSS in the offline phase and provides accurate positioning results in the online phase. This ultimately improves the efficiency, accuracy, and cost-effectiveness of radio map construction.
III-B Super-resolution
While these advancements are promising, there are still significant challenges to address in RF sensing. In particular, in many IoT deployments, the number of sensors is limited due to cost constraints, power requirements, or physical space limitations. This can result in large areas with minimal sensor coverage, leading to extremely sparse data. When there are large missing regions (e.g., 90% missing), GenAI models are more effective due to their ability to generate data that is not directly present in the input, enabling to infer missing data under large missing rates and complicated loss patterns [7]. Among them, DMs show its biggest potential for super-resolution data generation from extremely sparse latent spaces. DMs generate images through a process of gradually reducing noise, modeling the distribution of possible images at each step. This makes them effective at dealing with incomplete input. Despite the efficiency of DMs, they still bring the problem of low authenticity. This problem could be addressed by the idea of guided diffusion models, which utilize additional information, such as labels, text descriptions, or other types of metadata to guide the model for a more expected output. Building on this concept, LLMs can be employed to create customized guidance for a wide range of sensing tasks, such as tailoring the guiding prompts for radio map or any other patterns.
III-C Synthetic data generation
In IoT scenarios, data collection is challenging due to logistical, financial, and privacy constraints, as well as the variability and complexity of real-world environments. By leveraging AI models to generate high-fidelity synthetic data, it is possible to augment limited real-world data, enhancing the training and performance of RF signal processing algorithms. GenAI can improve model robustness by generating synthetic data from learned distributions, introducing controlled variability, and helping models generalize to unseen conditions. Specifically, GenAI can create data tailored to specific scenarios, such as varying environmental conditions, interference patterns, and device configurations, simulating diverse settings like urban environments, indoor spaces, and rural areas. In [8], a generative adversarial networks for RSS data augmentation is proposed to generate synthetic RSS data based on a small set of real collected labeled data, which achieved to efficiency improving localization accuracy. The inherent statistical character and use case examples in various domains illustrate that GenAI play a crucial role in addressing the data collection issues by generating synthetic data that enhances the training and performance of RF sensing models in IoT scenarios.
Figure 2: Illustration of GenAI techniques addressing challenges in uni-modal RF sensing for IoT systems through three key examples: completing missing RF sensor data, reconstructing radio maps from extremely sparse observations, and generating synthetic data to enhance model robustness and generalization across diverse environmental conditions.
IV GenAI empowered multi-modality RF sensing in IoT
In this section, we introduce the multi-modality RF sensing techniques in two parts, namely cross-modality RF sensing and multi-modality fusion for RF sensing (see Fig. 3). This division is motivated by the capabilities of GenAI to enhance RF sensing performance through distinct approaches. Cross-modality RF sensing leverages GenAI to correlate RF signals with another modality, such as images, to mitigate challenges like sparsity, interference and missing data. Meanwhile, multi-modality fusion combines data from multiple modalities using GenAI to create a comprehensive representation, overcoming the limitations of individual modalities and improving overall sensing performance.
To better highlight the potential of GenAI in each category of techniques, we provide the following definitions:
Definition 1
Cross-Modality RF sensing focuses on learning and utilizing the relationships between different modalities to infer one modality from another. It aims to handle scenarios where one modality may be missing or incomplete, by leveraging information from another modality.
Definition 2
Multi-modality fusion for RF sensing involves the simultaneous integration of multiple data modalities (e.g., visible images, LiDAR, audios), with RF signals, to create a unified representation that leverages the strengths of each modality and improve overall sensing accuracy and effectiveness. In this approach, it is assumed that multiple modalities are available, and it focuses on combining them to improve performance.
IV-A Cross-modal RF sensing with GenAI
In RF sensing scenarios, during the training phase, models may be trained using simulated RF data or data collected under controlled conditions where signals are easy to observe and interference are controlled. However, in real-world deployment, the actual RF signals might be sparse, noisy, or even unavailable due to environmental constraints or hardware limitations. GenAI-empowered cross-modal RF sensing techniques address the problems in various methodologies.
LLMs provide great potential for cross-modal learning with their strong capacity to enhance the generalizability of multimodal systems. A first strategy is to train universal models on vast datasets across numerous modalities, to encode these modalities into embeddings. More specifically, multi-modal LLMs can help to generate embeddings that effectively represent the underlying information from different modalities, therefor helping to enhance the generalizability of the universal models. The second direction, From a more precise perspective with LLM models, is to convert every modality into text modality and build the model based on text prompts. It has the potential to provide robust system for cross-modal sensing when one of the modalities is missing, as text can serve as a unified semantic space, leveraging the extensive zero-shot prediction capabilities of LLMs. For instance, [9] leverages LLMs and foundational models to create a unified semantic space using text. This approach shows the capacity of LLMs to enhance the generalizability of multimodal systems, allowing them to perform zero-shot predictions across different modality combinations during testing.
Figure 3: This figure illustrates the scenario of Generative AI-empowered multi-modal RF sensing applications. Three parts of different techniques are presented, namely the cross-modal generation, multi-modal fusion and LLM-enhanced interaction among sensors. The utilization of GenAI for multi-modal sensing makes the Internet of Sensors system more adaptive to dynamic environments and significantly improving overall efficiency.
VAEs provide great potential to learn to encode RF signals into a latent space and decode them into another modality. More specifically, based on the probabilistic nature of VAEs, they have the capacity to adopts a hierarchical approach to encode data into a distribution over latent variables, and generate diverse outputs in cross-modal tasks. VAEs can learn both joint and modality-specific distributions for different data types (e.g., RF signals and images), enabling effective cross-modality inference. This allows the model to leverage the strengths of each modality and make accurate predictions even when some data is missing or incomplete, by inferring the missing parts from the shared latent space. For instance, [10] introduced a Multimodal Hierarchical Variational Auto-encoder (MHVAE), where the single modality nature of VAEs is extended to a multimodal hierarchical setting, enabling the learning of modality-specific distributions and a joint-modality distribution for cross-modality inference. It showcasts the effectiveness of VAEs in learning rich multimodal representation and cross-modal generation.
DMs also offers a effective solution for synthesizing cross-modality data by leveraging existing datasets. Based on hierarchical statistical structure and iterative denoising process, DMs show its capacity for tasks requiring high-quality and detailed cross-modal outputs, such as text-to-image generation. What’s more, DMs can effectively capture and model complex relationships between different modalities, making them suitable for sophisticated cross-modal tasks, such as text-to-image generation. For instance, in [11], an effective method for synthesizing RF sensing data using cross-modal diffusion models to enhance the generalization capability of millimeter-wave (mmWave) sensing systems is proposed, demonstrating a remarkable enhancement in sensing and generalization capabilities across two distinct sensing tasks.
Besides, GenAI has the potential to revolutionize RF sensing by enabling personalized and context-aware applications that enhance user experiences and improve the efficiency of various systems. For instance, conditional GenAI techniques can be trained to generate data conditioned on specific inputs, such as text descriptions and contextual data like location or user preferences, therefor integrating contextual information to generate relevant and personalized outputs.
IV-B Multi-modal fusion with GenAI
The growing demand for real-world applications highlights the limitations of uni-modal sensing, especially for comprehensive environmental understanding. Integrating multi-modal sensing techniques can address this challenge by combining different data types (radio, optical, audio, thermal) with RF signals. This integration provides complementary information, enhancing the accuracy and robustness of sensing tasks. The redundancy from different sensors also improves the reliability and fault tolerance of wireless systems. For instance, in wireless localization, multi-modal techniques enhance positioning accuracy by providing redundant and complementary data, such as visible images, thermal images, and audio, which support and improve conventional uni-modal fingerprinting methods.
The advantages of multi-modal sensing come with new challenges, particularly the unique data patterns of each modality, making it hard to adapt models across modalities. Multi-modal fusion techniques, such as the Meta-Transformer [12], address this issue. A Meta-Transformer uses a data-to-sequence tokenizer to project data into a shared embedding space, a modality-agnostic encoder to process these embeddings, and task-specific heads for downstream predictions. This approach consolidates latent patterns for tasks like wireless positioning, autonomous driving, smart home technologies, and industrial automation.
Integrating GenAI, particularly LLMs, into multi-modality fusion offers transformative potential for RF sensing tasks. GenAI models are highly effective in understanding, generating, and translating complex data across various modalities. LLMs, with their advanced natural language processing, can integrate language as a supplementary modality alongside RF signals, images, and audio. This is particularly useful when sensor data is incomplete or sparse. For example, supplementary sensors can capture and transmit textual descriptions of the environment, such as object locations or specific conditions, with minimal data overhead, reducing costs. LLMs can then fuse this language data with RF signals and other modalities, enriching contextual understanding and filling gaps left by incomplete RF data, thereby enhancing overall sensing capabilities.
Figure 4: Illustration of the RF-to-image Generative AI model utilizing a text-prompt conditioned Transformer, and the reconstructed building obstacles image from real measurement. Initially, a Convolutional Neural Network (CNN) generates rough images from extremely sparse receive power samples of RF signals, which are blurred and noisy. These blurred images are then refined by GPT4o with a text prompt, enhancing image quality by removing noise and sharpening details. The figure highlights the significant improvements achieved through generative AI, demonstrating its capability to generate images with low distortion and high perceptual quality, thereby showcasing its advantages in RF sensing applications.
Enhancing RF Sensing with Language Descriptions: 1) Consider a scenario where a network of sensors is deployed in an environment for monitoring purposes. While the primary RF sensors capture signals that provide spatial and movement RF data, supplementary sensors record descriptive information about the environment. These descriptions might include details such as ”three people standing near the entrance,” ”a moving vehicle is approaching from the west,” or ”there is an obstruction in the path.” Transmitting these descriptions instead of raw data minimizes the bandwidth requirements and simplifies data collection. 2) Using LLMs, these language descriptions can be fused with the RF signal data to enhance downstream tasks such as localization, object detection, and activity recognition. LLMs can process and integrate the descriptive information, correlating it with the RF data to create a more complete and accurate representation of the monitored environment. Moreover, the flexibility of LLMs allows them to interpret the descriptive data, making them adapted to scenarios where traditional sensing methods might struggle. For instance, in dynamic environments where RF signals might be highly variable, the additional context provided by language descriptions can significantly improve the reliability and accuracy of the sensing system. A hybrid End-to-End learning framework for autonomous driving by combining basic driving imitation learning with LLMs based on multi-modality prompt tokens is proposed in [13], where end-to-end integration of visual and LiDAR sensory are input into learnable multi-modality tokens, and a hybrid setting is explored which use LLMs to help the driving model correct mistakes and complicated scenarios.
IV-C Case study: GenAI models for RF imaging
GenAI models have been widely used to generate data cross different modalities. For example, DALLE-2 uses diffusion model or auto-regressive model to generate image embedding from text embedding, plus contrastive learning to align the embeddings on pair-wise data. Moreover, causal transformer and U-Net are shown as effective architecture to implement the diffusion prior and diffusion generator. Inspired by this, we propose a RF to image GenAI model based text-prompt conditioned Transformer model. As illustrated in 4, the figure depicts the structure of our proposed method. Initially, a Convolutional Neural Network (CNN) is trained to generate images from extremely sparse receive power samples of RF signals, producing a rough representation often characterized by blur and noise. These blurred images are then fed into a Transformer model conditioned by a text prompt, which further enhances the image quality by leveraging its learned data distribution to remove noise and sharpen details. The comparison between the intermediate blurred images and the final output highlights the significant improvements achieved through generative AI. While traditional neural networks struggle to generate clear images from sparse data due to their limited capacity to learn exact data distributions, the Transformer model outperforms by providing high-quality reconstructions. This case study can show the capability of GenAI to generate images with both low distortion and high perceptual quality, highlighting its significant advantage in RF sensing applications.
V A foundation model for RF sensing and communication
Figure 5: GenAI model for multi-task wireless sensing. A unified GenAI model trained on generating visual or radio images is proposed and shown to integrate an LLM fine-tuned for downstream sensing and communication tasks.
While we have discussed the potential of GenAI models to enhance wireless sensing, an additional research question arises: can we pretrain a foundation GenAI model on sensory data that can be used for different sensing tasks and scenarios?
Inspired by the Meta-Transformer [12], a task-agnostic model trained for generating visual data, e.g., 3D environment image, can be used for computer vision tasks such as object detection, visualization and tracking. Similarly, a model trained for generating RF signals, e.g., 4D radio map, can support various wireless communication tasks, such as beam selection, mobility and spectrum management (Fig. 5), via task-specific fine-tuning. Specifically, cross-modality encoders can be trained to encode the raw data on a common latent space. This can be achieved by contrastive learning, following the CLIP [14] method to maximize cross-modality similarity between environment image and RF signal. Thereafter, a unified GenAI model is trained on the data embedding to generate the visual or radio environment representation, which can be realized by AR or DM trained on self-supervised learning (i.e. masked or next image/RF token prediction). Finally, the generated visual or RF embedding can be used by a multi-modal LLM to perform multiple downstream tasks in wireless sensing and communications. For example, a visual language model such as [15] can produce location of a targeted user from the constructed environment image, to complete the localization task. Similarly, a RF-language model can generate policies of optimal beams for tracking a user, to perform the beam management task. After pre-training on large wireless sensing datasets, we anticipate LLMs with embedded visual and radio data can perform multiple tasks in sensing and communications based on the user’s prompts. In doing so, we can build a unified model framework for sensing which does not require task specific retraining or fine-tuning, as shown by Fig. 5.
VI Challenges
In this section, we elaborate on several challenges of using generative AI encountered in RF sensing applications within IoT systems.
VI-A Prior information combination
One critical challenge lies in the effective utilization of prior information to enhance the accuracy of generated samples from sparse observations. In RF sensing, additional prior information, such as spatial configurations, historical signal patterns, and environmental characteristics, is often available and can significantly improve reconstruction accuracy by providing context and reducing ambiguity in the generated data. To leverage this, it is essential to integrate the prior information seamlessly with the architecture design in generative models, such as GANs or diffusion models, possibly by conditioning the generative model based on the prior information.
VI-B Distortion-perception tradeoff
GenAI techniques are highly effective in tasks requiring human or machine perception but encounter challenges when distortion is more critical, such as in radio map reconstruction. Therefore one challenge is to develop models that optimize both distortion and perception. One possible way to address this challenge is to integrate traditional neural networks, trained to minimize distortion, with GenAI models when dealing with problems that involve both distortion and perception.
VI-C Implementation of large models
Another challenge is the storage and computing of large GenAI models, particularly LLMs. These models enable intelligent edges in IoT systems but present storage and computing challenges on user equipment or edge devices. Efficient model compression techniques, such as pruning, quantization, or more developed techniques such as goal-oriented compression, are essential to deploy large models in resource-constrained equipment. Overcoming this challenge is essential for the practical deployment of GenAI-empowered RF sensing techniques.
VI-D Balancing synthetic and real data
While GenAI offers significant potential to enhance the training of neural networks by generating large amounts of synthetic data, it is crucial to carefully consider the balance between synthetic and real-world data. A major concern in the context of data augmentation with GenAI is determining the optimal amount of synthetic data to use in training a neural network. An insufficient amount of synthetic data can lead to inadequate training and poor generalization, while an excessive amount of synthetic data may introduce biases and reduce the model’s performance on real-world data.
VII Conclusions
In this paper, we explore the potential of GenAI for RF sensing to address data acquisition and transmission resource scarcity driven by the increasing demand for IoT systems. Our approach utilizes GenAI for efficient RF signal processing, incorporating uni-modal processing, cross-modal generation, and multi-modal fusion to effectively perform tasks such as radio map reconstruction and localization. We not only present these visions but also demonstrate them through practical case studies, discussing the challenges and opportunities, and proposing GenAI-based solutions as promising avenues for future developments in RF sensing. We foresee GenAI significantly benefiting various IoT applications: enhancing smart city operations by accurately predicting vehicle movements and reducing congestion through real-time data analysis; improving wearable IoT devices for better patient outcomes and reduced healthcare costs; and advancing predictive maintenance in industrial IoT by analyzing sensor data to predict equipment failures. GenAI has the potential to revolutionize our interactions with interconnected systems, leading to smarter, more efficient, and responsive IoT environments.